If log10X is normally distributed with mean 4 and variance 4, find the probability of 1.202<X<83180000.(Given that log10 1202=3.08, log10 8310=3.92 )
1
Expert's answer
2020-04-20T04:37:35-0400

Let require probability, "P=P( 1.202<X<83180000)"


"P( 1.202<X<83180000)" is convert to "\log_{10}X".

Then ,

probability"(P)" is "P( \log_{10}{1.202}<\log_{10}X<\log_{10}83180000)"


"\log_{10}{1.202}=\log_{10}{\frac{1202}{1000}}=\log_{10}{1202}-\log_{10}{1000}\\"

"=3.08-3=0.08"

"\log_{10}{83180000}=\log_{10}{8318*10^4}"

"=\log_{10}{8318}+\log_{10}{10000}\\"

"=3.92+4"

"=7.92"


"P=P(0.08<\log_{10}X<7.92)"

Then normalize the values,

"z=\frac{x-mean}{standard\ deviation}\\
z_1=\frac{0.08-4}{2}=-1.96\\
z_2=\frac{7.92-4}{2}=1.96\\"

"P=P(-1.96<z<1.96)\\
P=P(-1.96<z<0)+P(0<z<1.96)\\
P=2*P(0<z<1.96)\\"

using calculator or table,

"P=2*0.475\\
\bold{P=0.95}"